OpenStudy (anonymous):

In a mix of red and white pebbles in a bag in the ratio of 3 to 7. Find the probability that if 3 pebbles are chosen from the mixture at least one is white.

OpenStudy (confusionist):

43% 3/7= .4285714 Which simplifies to... 43%

OpenStudy (kc_kennylau):

P(At least one is white) =1-P(None is white) =1-(3/10)^3 =0.973

OpenStudy (confusionist):

Yes, that's a better explanation. I didn't finish mine. Sorry. ^

OpenStudy (anonymous):

@Confusionist I don't understand where you are intending to go with your answer @kc_kennylau 's answer is correct IF you are putting the pebbles back after each draw however if you are NOT putting the pebbles back, P(none is white) will no longer be 3/10 multiplied by itself three times. It will be 3/10 * 2/9 * 1/8. This reflects the diminishing pool of pebbles on each draw. P(at least one is white) = 1 - P(none is white) = 1 - (3/10)*(2/9)*(1/8) = 1 - 1/120 = 119/120 which is approximately 99.2%

OpenStudy (confusionist):

I misread the question, that's all. :) Wasn't intending to go anywhere except to go get a coffee so I can provide decent answers again instead of these tired four-in-the-morning ones. :)

OpenStudy (anonymous):

rock on :)

OpenStudy (kc_kennylau):

@campbell_st узр, sо fгзакιиg sιgиιfιсаиτ xD